OpenStudy (anonymous):

Suppose you are traveling straight from Los Angeles, California to Albuquerque, New Mexico, stopping overnight in Flagstaff, Arizona. The distance from L.A. to Albuquerque is 810 miles, and the road connecting the three cities is a straight line. If the distance from L.A. to Flagstaff is 120 miles greater than the distance from Flagstaff to Albuquerque, how far is L.A. from Flagstaff, and how far is Flagstaff from Albuquerque?

OpenStudy (anonymous):

come up with a basic equation such as: 810 = x + y where x is from LA to flagstaff and y is flagstaff to Albuquerque. Now you have the fact that FS to ALB is 120 miles greater so change the equation like this: ( known as substitution) 810 = (120 + y) + y 810 = 120 + 2y 690 = 2y y = 345 Now put this into the second equation we came up with. 810 = (120 + 345) + 345 x y That should be it. :)

OpenStudy (anonymous):

I menat LA to FS not FS to ALB sorry

OpenStudy (anonymous):

Let x be the distance between Flagstaff andAlbuquerque. Solve the following for x: (120 + x) + x = 810 x = 345 miles LA to Flagstaff is (120 + x) = (120+345) = 465 miles.